Find the radius of convergence for the series:
∑_(k=1)^∞▒〖-5x〗^k/(k+2)
1
Expert's answer
2020-11-08T18:01:04-0500

Use ratio test as,


"L""=\lim\nolimits_x\rarr\infin∣ \frac{a_{k+1}}{a_{k}}∣"


"=\lim\nolimits_x\rarr\infin∣ \frac{(-5x)^{k+1}}{(k+1)+2}\frac{k+2}{(-5x)^{k}}∣"


"=\lim\nolimits_x\rarr\infin∣ \frac{(-5x)^{k+1}}{(-5x)^{k}}\frac{k+2}{k+3}∣"


"=\lim\nolimits_x\rarr\infin∣ (-5x)\frac{k(1+\frac{2}{k})}{k(1+\frac{3}{k})}∣"


"=∣ (-5x)\frac{(1+0)}{(1+0)}∣"


"=\u2223 (-5x)\u2223"


"=\u2223 5x\u2223"



For convergence "\u2223 5x\u2223<1" , so


"-1<5x<1"


"-\frac{1}{5}<x<\frac{1}{5}"


The radius of convergence of the series is,


"R=\frac{1}{2}[\frac{1}{5}-(-\frac{1}{5})]"


"=\frac{1}{2}[\frac{1}{5}+\frac{1}{5}]"


"=\frac{1}{2}[\frac{2}{5}]"


"=\frac{1}{5}"


Therefore, the radius of convergence is "R=\frac{1}{5}"